### v1

@@ 4.1 Dispersive representation of nucleon gravitational form factors
$\operatorname{Disc}\left\langle N(p^{\prime})\bar{N}\left(p\right)\left|\hat{T}^{\mu\nu}(0)\right|0\right\rangle\propto\sum_{n}\left\langle\left.N(p^{\prime})\bar{N}\left(p\right)\right|n\right\rangle\left\langle n\left|\hat{T}^{\mu\nu}(0)\right|0\right\rangle^{*}\delta^{4}(p+p^{\prime}-p_{n})\ ,$ (72)
@@
with the trace and traceless parts given by [71]
@@
Once the spectral functions of the nucleon GFFs are determined, the DRs can be immediately formulated for the GFFs. In the timelike region, unitarity constraint implies that any FFs must vanish at infinite momentum transfer $t\rightarrow+\infty$ [130]. The exact behavior of the decaying power could be determined by the leading order pQCD [131, 132] and the Phragmèn-Lindelöf theorem [124, 133], which states that the spacelike asymptotic behavior of any FFs can be extended to any direction in the complex $t$ plane. Therefore, the GFFs satisfy the unsubtracted DR (for simplicity, we omit the superscript “$s$” or “$N$” for the nucleon GFFs in the following):
@@
However, these sum rules converge slowly. This phenomenon was discovered long ago in the dispersive analysis of the nucleon electromagnetic FFs [134, 135, 136, 137]. Here, following the methodology in Refs. [134, 135, 136, 137], we include additional effective zero-width poles with masses $m_{S,D}$ in the spectral functions, to simulate the contributions from possible high-lying excited meson states [71],
@@
where $F\in\{A,J,\Theta\}$, and the subscripts “$S$”, “$D$” denote $S$- and $D$-waves, respectively. The couplings $c_{S,D}$ are used to enforce the sum rules to be fulfilled. This formulation also allows for uncertainty estimates for the spectral densities by varying the effective pole masses $m_{S,D}$. We take the $m_{D}=1.8^{+0.4}_{-0.3}~{}$eV, which covers the mass range of $f_{2}(1565),f_{2}(1950)$ and $f_{2}(2010)$, and $m_{S}=1.6^{+0.2}_{-0.1}~{}$eV, covering the scalar resonances $f_{0}(1500)$ and $f_{0}(1710)$.
@@
From the $D$-term FF in Fig. 8, one gets the $D$-term as [71]
@@
which satisfies the positivity bound [149], $D\leq-0.20(2)$, and is in good agreement with the recent LQCD results, $-3.87(97)$ from the dipole fit and $-3.35(58)$ from the $z$-expansion fit in Ref. [70].
@@ 4.2 Spatial density profiles and mean squre radii
A stable numerical Fourier transformation requires knowledge of the complete $Q^{2}$-dependence of FFs, where $Q^{2}=-t$. The large $Q^{2}$ behavior of the spacelike nucleon GFFs is determined by the QCD hard scattering mechanism, up to logarithmic corrections, in Refs. [170, 131, 132, 38] to be $A(Q^{2})\sim J(Q^{2})\sim\Theta(Q^{2})\sim 1/Q^{4}$ and $D(Q^{2})\sim 1/Q^{6}$. With the above asymptotic behavior of the GFFs, Eq. (86) implies the superconvergence relations:
$\frac{1}{\pi}\int_{t_{\pi}}^{\infty}\mathrm{d}t^{\prime}\operatorname{Im}(A,J,\Theta)\left(t^{\prime}\right)=0\ .$ (98)
Unfortunately, numerical values of the left-hand side of the above relations highly depend on the high-energy tails of the spectral functions, which are poorly known and present significant challenges for an accurate description. Therefore, we adopt a $z$-expansion [171] method to extrapolate the GFFs in the spacelike region (for more details, see Appendix A). So far, all currently known calculations of spatial densities corresponding to GFFs, in particular at short distances, are more or less model-dependent, as we lack model-independent GFFs spanning in the whole range of $t$ from $t=0$ to $\infty$. Here, we use the central values of the GFFs derived in Ref. [71], as reviewed in Sec. 4.1, as inputs. These distributions are shown in Figs. 9 and 10.
@@
The 3D BF densities yield the mass, scalar trace, angular momentum and mechanical root-mean-square (rms) radii, $r_{i}\equiv\sqrt{\left\langle r_{i}^{2}\right\rangle}$ ($i=$Mass, $\Theta,J$, Mech), of the proton [3] as [71]: [^12] Here the “mass radius” $r_{\text{Mass}}$ is defined as the radius derived from the energy density; see, e.g., Ref. [172]. The radius of the trace density distribution is also called a “scalar radius” in the literature.
@@
$\sigma_{r}(b)=$ $\ m_{N}\int\frac{\mathrm{d}^{2}\bm{\Delta}_{\perp}}{(2\pi)^{2}}e^{-i\bm{b}_{\perp}\cdot\bm{\Delta}_{\perp}}\left[-\frac{1}{2b^{2}}\frac{1}{m_{N}^{2}}\frac{\mathrm{d}}{\mathrm{d}t}t^{\frac{3}{2}}D(t)\right]\equiv\sigma(r)+\frac{1}{2}\Pi(r)\ ,$ (110)
$\sigma_{t}(b)\equiv$ $\ \sigma(r)-\frac{1}{2}\Pi(r)\ ,$ (111)
$\sigma(r)=$ $\ \frac{1}{8m_{N}}\frac{1}{b}\frac{\mathrm{d}}{\mathrm{d}b}b\frac{\mathrm{d}}{\mathrm{d}r}\overline{D}(b)\ ,\quad\Pi(r)=-\frac{1}{4m_{N}}b\frac{\mathrm{d}}{\mathrm{d}r}\frac{1}{b}\frac{\mathrm{d}}{\mathrm{d}r}\overline{D}(b)\ ,$
where $b=|\bm{b}_{\perp}|$ and $\overline{D}(b)=\int\frac{\mathrm{d}^{2}\bm{\Delta}_{\perp}}{(2\pi)^{2}}e^{-i\bm{b}_{\perp}\cdot\bm{\Delta}_{\perp}}D\left(t\right)$. In Figs. 1112 and 13, we plot the above-mentioned 2D instant EF distributions as functions of the impact parameter $b$.
@@ 4.3.1 Matching at the next-to-leading order
$\mathcal{S}_{\pi\mathrm{N}}^{(2)}=\int\mathrm{d}^{4}x\sqrt{|g|}\left\{\frac{c_{8}}{8}R\bar{\psi}\psi+\frac{2c_{9}}{m_{N}}R^{\mu\nu}\left(\bar{\psi}e_{\mu}^{a}\gamma_{a}\nabla_{\nu}\psi-\nabla_{\nu}\bar{\psi}e_{\mu}^{a}\gamma_{a}\psi\right)\right\}.$ (114)
@@ 4.3.2 Matching at the next-to-next-to-leading order
In Fig. 14, we compare the predictions from NNLO BChPT to the dispersive results of the nucleon GFFs up to $-t=0.5~{}\mathrm{GeV}^{2}$. The NNLO BChPT predictions for $A(t)$ and $J(t)$ are in better agreement with the dispersive results than those for $\Theta(t)$ and $D(t)$. This can be understood as follows: while both $A(t)$ and $J(t)$ depend only on the $D$-wave $\pi\pi$, $\Theta(t)$ and $D(t)$ also depend on the $S$-wave $\pi\pi$; the contribution from the $\sigma$ meson in the $S$-wave $\pi\pi$ scattering amplitude, which has a much smaller mass than the $f_{2}(1270)$ resonance, is not effectively captured in the NNLO BChPT prediction but is contained in the DR analysis.
@@ 5 Summary and outlook
Appendix A $z$-expansion extrapolation of nucleon form factors
@@
$F^{N}\!\left(Q^{2}\right)=\mathcal{F}^{N}(z)=\sum_{k=0}^{k_{\max}}a_{k}z^{k}\ ,\quad z\left(t,t_{\mathrm{cut}},t_{0}\right)=\frac{\sqrt{t_{\mathrm{cut}}-t}-\sqrt{t_{\mathrm{cut}}-t_{0}}}{\sqrt{t_{\mathrm{cut}}-t}+\sqrt{t_{\mathrm{cut}}-t_{0}}}\ ,$ (124)
@@
Using the definition in Eq. (124), in the limit $Q^{2}\rightarrow\infty$, we have $z\rightarrow 1$ and $1-z\to 1/Q$, and the FF becomes
$\mathcal{F}^{N}(z)=\mathcal{F}^{N}(1)+\left.\sum_{n=1}^{k_{\max}}\frac{\mathrm{d}^{n}\mathcal{F}^{N}}{\mathrm{d}z^{n}}\right|_{z=1}(1-z)^{n}\sim\mathcal{F}^{N}(1)+\left.\sum_{n=1}^{k_{\max}}\frac{\mathrm{d}^{n}\mathcal{F}^{N}}{\mathrm{d}z^{n}}\right|_{z=1}\left(\frac{1}{Q^{n}}\right).$ (125)
If we assume that the FFs fall as positive powers of $1/Q$ at large $Q^{2}$, then we can write $\mathcal{F}^{N}(z\rightarrow 1)\propto 1/Q^{\ell}$, where $\ell$ is an integer. Combining this with Eq. (125), we obtain the constraints for the coefficients $a_{k}$:
$\mathcal{F}^{N}(1)=$ $\sum_{k=0}^{k_{\max}}a_{k}=0\ ,$ (126)
$\left.\frac{\mathrm{d}^{n}\mathcal{F}^{N}}{\mathrm{d}z^{n}}\right|_{z=1}=$ $\sum_{k=n}^{k_{\max}}\frac{k!}{(k-n)!}a_{k}=0,\quad n=1,\ldots,\ell-1\ .$ (127)
In this work, we take $\ell=4$ for $A,J,\Theta$ and $\ell=6$ for $D$ [131, 132]. We also set $k_{\max}=6$ and $k_{\max}=8$ for $A,J,\Theta$ and $D$, respectively, using $\left\{a_{0},a_{1},a_{2}\right\}$ as free parameters and leaving the rest of the coefficients $\left\{a_{3},\ldots,a_{6}\right\}$ and $\left\{a_{3},\ldots,a_{8}\right\}$ constrained.

### v2

@@ 4.1 Dispersive representation of nucleon gravitational form factors
[…]
@@
with the trace and traceless parts given by [73, 56]
@@
Once the spectral functions of the nucleon GFFs are determined, the DRs can be immediately formulated for the GFFs. In the timelike region, unitarity constraint implies that any FFs must vanish at infinite momentum transfer $t\rightarrow+\infty$ [137]. The exact behavior of the decaying power could be determined by the leading order pQCD [138, 139] and the Phragmèn-Lindelöf theorem [131, 140], which states that the spacelike asymptotic behavior of any FFs can be extended to any direction in the complex $t$ plane. Therefore, the GFFs satisfy the unsubtracted DR (for simplicity, we omit the superscript “ss” or “NN” for the nucleon GFFs in the following):
@@
However, these sum rules converge slowly, especially for $\Theta(t)$; see the weighted spectral functions for $(A,J,\Theta)(t)/t$ in Fig. 10. A detailed discussion of sum rule saturation for nucleon GFFs appears in Appendix A. This phenomenon was discovered long ago in the dispersive analysis of the nucleon electromagnetic FFs [141, 142, 143, 144]. Here, following the methodology in Refs. [141, 142, 143, 144], we include additional effective zero-width poles with masses $m_{S,D}$ in the spectral functions, to simulate the contributions from possible high-lying excited meson states [73],
@@
where $F\in\{A,J,\Theta\}$, and the subscripts “SS”, “DD” denote $S$- and $D$-waves, respectively. The couplings $c_{S,D}$ are used to enforce the sum rules to be fulfilled. This formulation also allows for uncertainty estimates for the spectral densities by varying the effective pole masses $m_{S,D}$. We take the $m_{D}=1.8^{+0.4}_{-0.3}$~GeV, which covers the mass range of $f_{2}(1565),f_{2}(1950)$ and $f_{2}(2010)$, and $m_{S}=1.6^{+0.2}_{-0.1}$~GeV, covering the scalar resonances $f_{0}(1500)$ and $f_{0}(1710)$.
Figure 10: Results of the nucleon weighted spectral functions for $A(t)/t,J(t)/t$ and $\Theta(t)/t$, respectively.
@@
From the $D$-term $F$ in Fig. 11, one gets the $D$-term as [73]
@@
which satisfies the positivity bound [156], $D\leq-0.20(2)$, and is in good agreement with the recent LQCD results, 3.87(97)-3.87(97) from the dipole fit and 3.35(58)-3.35(58) from the $z$-expansion fit in Ref. [72].
@@ 4.2 Spatial density profiles and mean squre radii
A stable numerical Fourier transformation requires knowledge of the complete $Q^{2}$-dependence of FFs, where $Q^{2}=-t$. The large $Q^{2}$ behavior of the spacelike nucleon GFFs is determined by the QCD hard scattering mechanism, up to logarithmic corrections, in Refs. [177, 138, 139, 39] to be $A(Q^{2})\sim J(Q^{2})\sim\Theta(Q^{2})\sim 1/Q^{4}$ and $D(Q^{2})\sim 1/Q^{6}$. With the above asymptotic behavior of the GFFs, Eq. (86) implies the superconvergence relations; see Appendix A. Unfortunately, numerical values of the left-hand side of the above relations highly depend on the high-energy tails of the spectral functions, which are poorly known and present significant challenges for an accurate description. Therefore, we adopt a $z$-expansion [178] method to extrapolate the GFFs in the spacelike region (for more details, see Appendix B). So far, all currently known calculations of spatial densities corresponding to GFFs, in particular at short distances, are more or less model-dependent, as we lack model-independent GFFs spanning in the whole range of $t$ from $t=0$ to $\infty$. Here, we use the central values of the GFFs derived in Ref. [73], as reviewed in Sec. 4.1, as inputs. These distributions are shown in Figs. 12 and 13.
@@
The 3D BF densities yield the mass, scalar trace, angular momentum and mechanical root-mean-square (rms) radii, $r_{i}\equiv\sqrt{\left\langle r_{i}^{2}\right\rangle}$ (i=i=Mass, $\Theta,J$, Mech), of the proton [3] as [73]: [^13] Here the “mass radius” $r_{\text{Mass}}$ is defined as the radius derived from the energy density; see, e.g., Ref. [179]. The radius of the trace density distribution is also called a “scalar radius” in the literature.
@@
$\sigma_{r}(b)=$ $\ m_{N}\int\frac{\mathrm{d}^{2}\bm{\Delta}_{\perp}}{(2\pi)^{2}}e^{-i\bm{b}_{\perp}\cdot\bm{\Delta}_{\perp}}\left[-\frac{1}{2b^{2}}\frac{1}{m_{N}^{2}}\frac{\mathrm{d}}{\mathrm{d}t}t^{\frac{3}{2}}D(t)\right]\equiv\sigma(b)+\frac{1}{2}\Pi(b)\ ,$ (109)
$\sigma_{t}(b)\equiv$ $\ \sigma(b)-\frac{1}{2}\Pi(b)\ ,$ (110)
$\sigma(b)=$ $\ \frac{1}{8m_{N}}\frac{1}{b}\frac{\mathrm{d}}{\mathrm{d}b}b\frac{\mathrm{d}}{\mathrm{d}b}\overline{D}(b)\ ,\quad\Pi(b)=-\frac{1}{4m_{N}}b\frac{\mathrm{d}}{\mathrm{d}b}\frac{1}{b}\frac{\mathrm{d}}{\mathrm{d}b}\overline{D}(b)\ ,$
where $b=|\bm{b}_{\perp}|$ and $\overline{D}(b)=\int\frac{\mathrm{d}^{2}\bm{\Delta}_{\perp}}{(2\pi)^{2}}e^{-i\bm{b}_{\perp}\cdot\bm{\Delta}_{\perp}}D\left(t\right)$. Note that the asymptotic large-$b$ behaviour can be determined by the infrared physics contained in unitarity relations [56].
In Figs. 1415 and 16, we plot the above-mentioned 2D instant EF distributions as functions of the impact parameter $b$.
@@ 4.3.1 Matching at the next-to-leading order
$\mathcal{S}_{\pi\mathrm{N}}^{(2)}=\int\mathrm{d}^{4}x\sqrt{|g|}\left\{\frac{c_{8}}{8}R\bar{\psi}\psi+\frac{ic_{9}}{m_{N}}R^{\mu\nu}\left(\bar{\psi}e_{\mu}^{a}\gamma_{a}\nabla_{\nu}\psi-\nabla_{\nu}\bar{\psi}e_{\mu}^{a}\gamma_{a}\psi\right)\right\}.$ (113)
@@ 4.3.2 Matching at the next-to-next-to-leading order
In Fig. 17, we compare the predictions from NNLO BChPT to the dispersive results of the nucleon GFFs up to $t=0.5~\mathrm{GeV}^{2}$. The NNLO BChPT predictions for $A(t)$ and $J(t)$ are in better agreement with the dispersive results than those for $\Theta(t)$ and $D(t)$. This can be understood as follows: while both $A(t)$ and $J(t)$ depend only on the $D$-wave $\pi\pi$, $\Theta(t)$ and $D(t)$ also depend on the $S$-wave $\pi\pi$; the contribution from the $\sigma$ meson in the $S$-wave $\pi\pi$ scattering amplitude, which has a much smaller mass than the $f_{2}(1270)$ resonance, is not effectively captured in the NNLO BChPT prediction but is contained in the DR analysis.
@@ 5 Summary and outlook
Appendix A Sum rules of gravitational form factors
Converting dispersion relations into sum rules relies on low- or high-energy expansions. Expanding both sides of the dispersion relations (86) in a Taylor series around $t\to\infty$ and equating the coefficients, one obtains the following superconvergence sum rules [56]:
$\frac{1}{\pi}\int_{t_{\pi}}^{\infty}\mathrm{d}t~t^{n-1}\operatorname{Im}\left(A,J,\Theta\right)(t)=0,\quad n=1,2,$ (A1)
that is, $\lim_{t\to\infty}t^{n}\times\left(A,J,\Theta\right)(t)=0$ due to the extra logarithmic suppression by the pQCD coupling [56]. If we have a set of perfect representations of GFFs at any energy, these sum rules must be satisfied. However, without sufficient high-energy inputs, the dispersive results are only applicable at relatively low energies, and we cannot expect these sum rules to be fully saturated by the purely dispersive components, especially for the superconvergence sum rules, which certainly require the contribution of highly excited state poles to cancel each other out. Table A1 shows the degree of saturation of each sum rule from the $\pi\pi,K\bar{K}$ continua. Here, two integration limits are presented. The first is up to $\sqrt{t}=1.3$~GeV, which corresponds to the contribution from the $\pi\pi,K\bar{K}$ continua that are precisely known; above this energy, the $S$-wave two-channel approximation starts to break down, and a simple extrapolation without any obvious resonant structure is adopted for the $S$-wave phase shifts. The second is at the two-nucleon threshold, $\sqrt{t}=2m_{N}$. One indeed finds that the superconvergence sum rules require more high energy inputs to be saturated.
Table A1: Degree to which the $\pi\pi,K\bar{K}$ continua saturate the nucleon GFF sum rules. Three GFFs, $(A,2J,\Theta/m_{N})$, are normalized to unity at $t=0$ for easy illustration. Here, $n$ is the order of the weight in the superconvergence sum rules in Eq. (A1), and the case of $n=0$ corresponds to Eq. (87). The upper limit of the integration in the sum rules is specified by $\Lambda^{2}$ with $\Lambda=1.3$ GeV or $\Lambda=2m_{N}$.
$n=0$ $n=1$ $n=2$
$\Lambda$ $1.3$ GeV $2m_{N}$ $1.3$ GeV $2m_{N}$ $1.3$ GeV $2m_{N}$
AA $1.35$ $1.77$ $1.79$ $2.58$ $2.49$ $4.04$
$2J$ $1.19$ $1.42$ $1.43$ $1.88$ $1.91$ $2.76$
$\Theta/m_{N}$ $1.32$ $2.60$ $0.67$ $3.71$ $0.59$ $8.09$
Analogous to the superconvergence sum rule of nucleon GFFs (A1), the meson GFFs $A^{\pi,K}$ also satisfy similar superconvergence sum rules [38]
$\frac{1}{\pi}\int_{t_{\pi}}^{\infty}\mathrm{d}t~\operatorname{Im}A^{\pi,K}(t)=0,$ (A2)
that is, $\lim_{t\to\infty}tA^{\pi,K}(t)=0$. The degrees of saturation of the above sum rules for the pion and kaon from the $\pi\pi,K\bar{K}$ continua are 1.86 and 0.77, respectively, when the integration upper limit is set to 2 GeV.
Appendix B $z$-expansion extrapolation of nucleon form factors
@@
$F^{N}\!\left(Q^{2}\right)=\mathcal{F}^{N}(z)=\sum_{k=0}^{k_{\max}}a_{k}z^{k}\ ,\quad z\left(t,t_{\mathrm{cut}},t_{0}\right)=\frac{\sqrt{t_{\mathrm{cut}}-t}-\sqrt{t_{\mathrm{cut}}-t_{0}}}{\sqrt{t_{\mathrm{cut}}-t}+\sqrt{t_{\mathrm{cut}}-t_{0}}}\ ,$ (B1)
@@
Using the definition in Eq. (B1), in the limit $Q^{2}\rightarrow\infty$, we have $z\rightarrow 1$ and $1-z\to 1/Q$, and the FF becomes
$\mathcal{F}^{N}(z)=\mathcal{F}^{N}(1)+\left.\sum_{n=1}^{k_{\max}}\frac{\mathrm{d}^{n}\mathcal{F}^{N}}{\mathrm{d}z^{n}}\right|_{z=1}(1-z)^{n}\sim\mathcal{F}^{N}(1)+\left.\sum_{n=1}^{k_{\max}}\frac{\mathrm{d}^{n}\mathcal{F}^{N}}{\mathrm{d}z^{n}}\right|_{z=1}\left(\frac{1}{Q^{n}}\right).$ (B2)
If we assume that the FFs fall as positive powers of $1/Q$ at large $Q^{2}$, then we can write $\mathcal{F}^{N}(z\rightarrow 1)\propto 1/Q^{\ell}$, where $\ell$ is an integer. Combining this with Eq. (B2), we obtain the constraints for the coefficients $a_{k}$:
$\mathcal{F}^{N}(1)=$ $\sum_{k=0}^{k_{\max}}a_{k}=0\ ,$ (B3)
$\left.\frac{\mathrm{d}^{n}\mathcal{F}^{N}}{\mathrm{d}z^{n}}\right|_{z=1}=$ $\sum_{k=n}^{k_{\max}}\frac{k!}{(k-n)!}a_{k}=0,\quad n=1,\ldots,\ell-1\ .$ (B4)
In this work, we take $\ell=4$ for $A,J,\Theta$ and $\ell=6$ for $D$ [138, 139]. We also set $k_{\max}=6$ and $k_{\max}=8$ for $A,J,\Theta$ and $D$, respectively, using $\left\{a_{0},a_{1},a_{2}\right\}$ as free parameters and leaving the rest of the coefficients $\left\{a_{3},\ldots,a_{6}\right\}$ and $\left\{a_{3},\ldots,a_{8}\right\}$ constrained. The corresponding fit parameters are listed in Table B1.
Table B1: Parameters in the $z$-expansion extrapolation, cf. Eq. (B1), obtained by fitting to the dispersive results of the four GFFs, $A,J,\Theta$ and $D$. The errors listed here are statistical only.
AA $J$ $\Theta/m_{N}$ $D$
$a_{0}$ $0.568(1)$ $0.264(0)$ $0.079(1)$ 1.080(1)-1.080(1)
$a_{1}$ 1.208(4)-1.208(4) 0.572(2)-0.572(2) 0.638(3)-0.638(3) $4.202(7)$
$a_{2}$ 0.120(8)-0.120(8) 0.054(4)-0.054(4) $1.426(9)$ 4.276(25)-4.276(25)